# Rethinking Efficient and Effective Point-based Networks for Event Camera Classification and Regression: EventMamba

### v1

@@ IV-F2 Point number
In contrast to previous work [8, 49], we increased the number of input points to 2048 on large action recognition datasets, and to determine this significant impact factor we verified it on HMDB51-DVS. The experiments compare the accuracy at training 100 epoch, EventMamba with 2048 points has 82.1% accuracy on the testset, with 1024 points has 52.5% accuracy on the testset, and with 512 points has 41.82% accuracy on the testset. This shows that increasing the number of input points can significantly increase the performance of the model on large action recognition datasets.
TABLE X: Abation of temporal modules on IJRR.
Network w/o Mamba & TAgg EventMamba
Parameter 0.245M 0.246M
FLOPs 0.392G 0.392G
shapes_rotation 0.007m,1.481 0.004m,1.091
box_translation 0.019m,1.237 0.016m,0.810
shapes_translation 0.015m,0.827 0.010m,0.600
dynamic_6dof 0.018m,1.154 0.014m,0.911
hdr_poster 0.029m,1.744 0.015m,0.985
poster_translation 0.021m,0.991 0.013m,0.587
Average 0.018m,1.239 0.012m,0.831(+33.1%)
V Conclusion
We present EventMamba, an efficient and effective Point Cloud network meticulously crafted for the classification and regression tasks associated with event cameras. Our evaluation encompasses seven predominant datasets, yielding outstanding outcomes. EventMamba not only demonstrates high accuracy but also proves to be apt for applications requiring low latency, even when computational resources are constrained.
In our future work, we will devote ourselves to the architecture development of the combination of Point Cloud processing network and spiking neural network (SNN) and can design a neuromorphic computing algorithm that is hardware-software friendly.

### v2

@@ 4.6.2 Point Number
In contrast to previous work [7, 55], we increased the number of input points to 2048 on large action recognition datasets, and to determine this significant impact factor, we verified it on HMDB51-DVS. The experiments compare the accuracy at training 100 epoch as described in Table XI, EventMamba with 2048 points has 82.1% accuracy on the testset, with 1024 points has 76.6% accuracy on the testset, and with 512 points has 67.9% accuracy on the testset. This shows that increasing the number of input points can significantly increase the performance of the model on large action recognition datasets.
4.6.3 Mamba or LSTM?
Table XI and XII show the performance of the two explicit feature extractors on the HMDB51-DVS and IJRR datasets, respectively. We try to ensure that both models consume an equitable computational resource, although the Mamba-based model still slightly surpasses the LSTM-based model. On the HMDB51-DVS dataset, Mamba demonstrated a notable performance advantage, achieving an accuracy of 82.1%, which exceeded LSTM’s accuracy of 70.8% by 16%. Likewise, on the IJRR dataset, Mamba exhibited superior performance with its average Mean Squared Error (MSE) outperforming LSTM’s by 8%. Additionally, the Mamba-based model’s inference is faster than LSTM, at 5.5ms compared to 6.8ms, respectively. Ignoring the slight increase in computational resources, we believe that Mamba is superior to LSTM.
5 Conclusion
In this paper, we rethink the Point Cloud representation to address event-based tasks by introducing a spatio-temporal hierarchical structure. This innovative paradigm effectively retains fine-grained temporal information between events and aligns with the inherently sparse and asynchronous characteristics of event camera. Comprehensive experiments substantiate the efficacy of this approach in both regression and classification tasks, and we will derive this to more event-based tasks in the future.